OpenStudy (anonymous):

With the restriction x ≠ 0, Which of the following is the simplified form of 6x^4/4x^9 * 12x^5/3x^2 6/x^2 6x^8 6/x^8 6x^2

OpenStudy (kinggeorge):

6/x^2 Because x^4/x^9 * x^5/x^2 = 1/x^2 and 6/4 * 12/3=6